OpenStudy (anonymous):

Write the equation of the line and draw the graph of the line through each pair of points given. Use either the slope-intercept equation or by using x- and y-intercepts >>(2, -1) and (0, 4)

OpenStudy (helder_edwin):

do you know this \[ m=\frac{y_0-y_1}{x_0-x_1} \] ?

OpenStudy (anonymous):

no..

OpenStudy (helder_edwin):

the equation gives you the slope of the line passing through 2 given points you have to points you need to find the slope. just plug in the number you are given!

OpenStudy (anonymous):

i dont know which ones to plug

OpenStudy (mathteacher1729):

@coolaidd -- do you have graph paper handy?

OpenStudy (anonymous):

i actually do not.. but i could probably write it on regular paper?

OpenStudy (mathteacher1729):

It's much easier with graph paper.

OpenStudy (anonymous):

i know..but i dont have it.

OpenStudy (mathteacher1729):

Do you have geogebra? You don't even need to install it. YOu can run it off the web.

OpenStudy (mathteacher1729):

The key here (I think) Is visually understanding what slope means, instead of just "plugging it into a formula".

OpenStudy (mathteacher1729):

Can you find : 1) rise 2) run 3) y-intercept By looking at the picture?

OpenStudy (anonymous):

3,2?

OpenStudy (mathteacher1729):

Well the line is sloping downward, so the "rise" is -3 not positive 3.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

so that is exactly how you plot it..what would be the equation?

OpenStudy (anonymous):

hello?

OpenStudy (mathteacher1729):

The slope-intercept equation will have the form y = mx + b Where m = slope (that's rise / run) b = the y intercept (the y-value where the line touches the y-axis) You literally pop those numbers in, and you're done. :)

OpenStudy (anonymous):

y=-3/2x+4 ?

OpenStudy (mathteacher1729):

Horay! :D

OpenStudy (anonymous):

am i right??

OpenStudy (anonymous):

lol id never thought id get it

OpenStudy (mathteacher1729):

You did! :D Seriously, use geogebra, it makes this sort of stuff very very easy to do. :) I can show you how to use if you'd like.

OpenStudy (anonymous):

thank youu soo much! when i get more graphing questions I will definitely message you! =)

OpenStudy (mathteacher1729):

Here are the instructions.

OpenStudy (anonymous):

thankss